# ReaCritic: Large Reasoning Transformer-based DRL Critic-model Scaling For Heterogeneous Networks

### v1

@@ III-A1 Data Rate
where $b_{{\rm{u}},m},b_{{\rm{d}},m}\in[0,1]$ are the normalized bandwidth allocation factors, and $B_{\mathrm{u},m}$ and $B_{\mathrm{d},m}$ are the maximum available bandwidths for uplink and downlink transmissions.
III-A2 Energy Efficiency
The uplink and downlink energy efficiencies, denoted by $\eta_{\mathrm{u},m}$ and $\eta_{\mathrm{d},m}$ (in bits per Joule), are defined as the ratio of achievable data rate to the corresponding transmission power:
@@ III-A2 Energy Efficiency
where the denominator represents the actual transmit power allocated to the $m_{\rm th}$ user for uplink and downlink, respectively.
III-A3 Service Latency
@@ III-B Problem Formulation
where $\alpha_{m}^{(i)}$, $i\in{1,\ldots,5}$, are user-specific preference weights corresponding to uplink data rate, downlink data rate, uplink energy efficiency, downlink energy efficiency, and service latency, respectively. The constraints jointly ensure that each user’s power, bandwidth, and compute allocations remain within feasible per-user limits, while also satisfying global capacity constraints across all users.
IV Large Reasoning Transformer-based DRL Critic-model Scaling
In this section, we formulate the Markov Decision Process (MDP) for the DRL-based HetNet management and introduce ReaCritic, a large reasoning transformer-based critic-model scaling scheme designed to address key challenges in HetNets.
IV-A DRL Markov Decision Process Formulation
DRL offers a principled framework for sequential decision-making by enabling an agent to interact with an environment $\mathcal{E}$. At each discrete time step $t$, the agent observes a state ${\bm{s}}_{t}\in\mathcal{S}$ from the state space and selects an action ${\bm{a}}_{t}\in\mathcal{A}$ according to its learned policy. The environment responds with a reward signal $r_{t}^{\mathcal{E}}\in\mathbb{R}$, and transitions to a new state ${\bm{s}}_{t+1}$ based on a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The learning objective is to optimize the expected cumulative reward by modeling the agent–environment interaction as a MDP.
@@ IV-A DRL Markov Decision Process Formulation
    where $\tau_{m}\in\{0,1,\ldots,\tau_{\max}\}$ denotes the user type. Each user type corresponds to a distinct set of preferences over multiple objectives, encoded by weights $\alpha_{m}^{(i)}$ $\left(i=1,\ldots,5\right)$ in the utility function $Q_{m}$. These preference profiles are hidden from the network manager, introducing uncertainty into the learning process and increasing the complexity of policy adaptation.
@@ IV-B ReaCritic Design
ReaCritic introduces a two-dimensional reasoning architecture designed to enhance the expressiveness and generalization capability of the critic network. The first stage, denoted as Horizontal Reasoning (HRea), performs token-level expansion by generating multiple parallel state-action embeddings to simulate diverse trajectories or perspectives. This enriches the input space with latent behavioral variations and reduces sensitivity to local minima. The second stage, denoted as Vertical Reasoning (VRea), applies stacked Transformer blocks to process these tokens hierarchically, enabling deep abstraction and structured value propagation through multi-head self-attention and feed-forward layers.
As illustrated in Fig. 3, the HRea and VRea steps together enable ReaCritic to support fine-grained relational modeling and robust Q-value estimation in complex, high-dimensional HetNets. The whole procedure is provided in Algorithm 1.
@@
Input:
  • Environment: Markov environment $\mathcal{E}$, replay buffer $\mathcal{D}$
  • DRL Algorithm $\mathcal{A}$: discount factor $\gamma$, actor network $\pi_{\theta}$ (if have), target networks $Q_{\bar{\phi}}$, $\pi_{\bar{\theta}}$, learning rates $\eta_{Q}$, $\eta_{\pi}$, mini-batch size $M$
  • ReaCritic: critic network $Q_{\phi}$ with parameters $\phi$, HRea steps $H$, VRea steps $V$
@@
6:         Sample action $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ according to DRL algorithm $\cal A$’s policy or other algorithm-specific rules
7:         Execute $a_{t}$, observe $r_{t}$, $s_{t+1}$, and termination flag $d_{t}\in\{0,1\}$
@@
11:     # Update phase
12:     Sample mini-batch $\{(s_{t},a_{t},r_{t},s_{t+1})\}$ from $\mathcal{D}$
13:     (ReaCritic Update)
14:     ## Horizontal Reasoning:
15:     Compute state-action embedding $\mathbf{z}_{0}$ using (18)
16:     Expand into $H$ tokens using (19) and add exploration noise via (20)
17:     ## Vertical Reasoning:
18:     Refine tokens through $V$ transformer blocks using (21)
19:     Aggregate token representations via step-wise attention (22)
20:     Predict final Q-value using (23)
21:     (Critic loss and Update)
22:     Compute target Q-value:
@@
23:     Compute critic loss: $\mathcal{L}_{Q}=\frac{1}{M}\sum_{i=1}^{M}(Q_{\phi}(s_{t},a_{t})-y_{t})^{2}$
24:     Update critic parameters: $\phi\leftarrow\phi-\eta_{Q}\nabla_{\phi}\mathcal{L}_{Q}$
25:     if actor $\pi_{\theta}$ exists then
26:         Compute policy gradient using (25)
27:         Update actor:  $\theta\leftarrow\theta+\eta_{\pi}\nabla_{\theta}J$
28:     end if
29:end for
Output: Trained DRL algorithm $\cal A$
IV-B1 Horizontal Reasoning Path
@@
However, unlike LLMs that sequentially generate tokens, DRL critic networks operate under a single-shot evaluation paradigm. Therefore, to mimic multi-step reasoning, ReaCritic introduces the concept of horizontal reasoning expansion: instead of processing a single embedding, the model replicates the input state-action token multiple times and augments each copy with a unique positional embedding, effectively simulating parallel hypothetical reasoning paths.
@@
where $\mathbf{W}_{\text{embed}}\in\mathbb{R}^{(d_{\bm{s}}+d_{\bm{a}})\times d_{\rm h}}$ is the trainable projection matrix, and $\mu(\cdot)$ denotes layer normalization. To enable multi-path reasoning, this base embedding is then expanded into a sequence of $H$ horizontally parallel reasoning tokens as
@@
This horizontally extended token sequence is then passed to the subsequent vertical reasoning transformer layers, where latent value estimation is refined through attention-based abstraction. The key intuition behind HRea is to provide the critic network with multiple perspectives on the same decision point, enhancing robustness, exploration, and its ability to model high-order interactions in complex environments.
IV-B2 Vertical Reasoning Path
@@
This design is motivated by the idea of latent multi-step reasoning in LLMs [39], where deep transformer layers repeatedly refine representations in hidden space to support complex, context-dependent computation. Here, VRea does not explicitly produce symbolic CoT, but instead performs recursive abstraction within the vector space. This allows the critic to simulate structured inference trajectories by reprocessing the latent input across multiple transformer layers, enabling deeper exploration and more robust Q value modeling.
VRea also shares conceptual links with diffusion policy [40] and iterative policy refinement, where each step in the stack can be interpreted as an implicit planning operation conditioned on both prior layers and current observations. Compared with simple feed-forward architectures, the VRea stack increases reasoning depth without increasing parameter count proportionally, and facilitates richer representation learning through hierarchical feature composition.
@@
where $\Psi^{(v)}(\cdot)$ denotes the $v_{\rm th}$ transformer block. Each layer refines the horizontal token sequence by integrating long-range dependencies and capturing high-order interactions through attention and non-linear transformation. After $V$ layers, the final representation $\mathbf{Z}^{(V)}\in\mathbb{R}^{B\times H\times d_{\rm h}}$ is aggregated using attention over the $H$ horizontal reasoning steps as
@@
This vertical reasoning mechanism enables ReaCritic to synthesize structured value estimates by recursively abstracting state-action pairs, resulting in improved generalization and temporal stability across complex HetNet scenarios.
IV-B3 Integration with DRL algorithms
Figure 4: Integration of ReaCritic into DRL algorithms. The part $1$ shows the reasoning-augmented SAC pipeline. Parts $2\sim 5$ illustrate compatibility with other actor-critic methods, including TD3, A3C, DDPG, and PPO.
As shown in Fig. 4, ReaCritic is designed as a plug-and-play critic module compatible with standard actor-critic DRL algorithms. While our implementation is based on the Soft Actor-Critic (SAC) framework [41], the design can be generalized to other value-based methods such as Twin Delayed Deep Deterministic (TD3) [42], Deep Deterministic Policy Gradient (DDPG) [43], and Asynchronous Advantage Actor-Critic (A3C) algorithms [44].
@@ IV-B3 Integration with DRL algorithms
IV-B4 Theoretical Complexity and Convergence Analysis
@@
The per-update computational complexity of ReaCritic is $\mathcal{O}(BH^{2}d_{\rm h}+BHVd_{\rm h}^{2})$, where $B$ is the batch size, $H$ is the number of HRea steps, $V$ is the number of vertical Transformer blocks, and $d_{\rm h}$ is the hidden dimension. The first term corresponds to the cost of multi-head self-attention across $H$ tokens per layer, while the second term captures the complexity of feed-forward transformations in each Transformer block. This structure scales linearly with $B$ and $V$, and quadratically with $H$ and $d_{\rm h}$. Compared to standard MLP-based critics with complexity $\mathcal{O}(Bd_{\rm h}^{2})$, the added cost remains moderate for lightweight settings with small $H$ and $V$, e.g., $H{=}4$, $V{=}3$, and can be efficiently batched on modern GPUs or edge accelerators. More importantly, ReaCritic is designed to support scaling along both the depth ($V$) and breadth ($H$) dimensions, enabling deployment of large reasoning-capable critics. This two-dimensional scaling path aligns with recent trends in inference-time compute augmentation and allows the model to serve both compact on-device learning and large-scale centralized training when compute resources permit.
@@
Regarding convergence, although ReaCritic increases model depth and nonlinearity, its use of pre-normalized residual blocks and standard Transformer modules supports stable training. Since it is integrated into conventional DRL update rules, e.g., Bellman backup and policy gradients, the theoretical convergence properties of the base algorithms, such as SAC or TD3, remain valid, assuming the critic maintains universal approximation capacity. The added reasoning depth improves the expressiveness of the value function, which can reduce approximation bias and accelerate convergence. Additionally, stepwise attention and horizontal perturbation act as implicit regularization, improving stability and credit assignment in complex, non-stationary environments.
V Numerical Results

### v2

@@ 3.1.1 Data Rate
where $b_{{\rm{u}},m},b_{{\rm{d}},m}\in[0,1]$ are the normalized bandwidth allocation factors, and $B_{\mathrm{u},m}$ and $B_{\mathrm{d},m}$ are the maximum bandwidths for uplink and downlink transmissions.
3.1.2 Energy Efficiency
The uplink and downlink energy efficiencies, denoted by $\eta_{\mathrm{u},m}$ and $\eta_{\mathrm{d},m}$ (in bits per Joule), are defined as the ratio of achievable data rate to transmission power:
@@ 3.1.2 Energy Efficiency
where the denominator represents the transmit power allocated to the $m_{\rm th}$ user for uplink and downlink, respectively.
3.1.3 Service Latency
@@ 3.2 Problem Formulation
where $\alpha_{m}^{(i)}$, $i\in{1,\ldots,5}$, are user-specific preference weights corresponding to uplink data rate, downlink data rate, uplink energy efficiency, downlink energy efficiency, and service latency, respectively. The specific structure of the power constraints reflects the fundamental difference between uplink and downlink hardware limitations. While downlink transmissions are typically subject to a total power budget at the base station, the uplink power allocation is dictated by the maximum transmit capability of individual user equipment. Therefore, the constraint on $p_{\mathrm{u},m}$ is applied element-wise to characterize the physical power amplifier limits of each mobile terminal. If a total uplink power budget across subcarriers is required in certain implementations, it can be incorporated by adding an additional summation constraint without affecting the proposed ReaCritic architecture. The constraints jointly ensure that each user’s power, bandwidth, and compute allocations remain within feasible per-user limits, while also satisfying global capacity constraints across all users.
4 Reasoning Transformer-based DRL Critic-model Scaling
In this section, we formulate the MDP for the DRL-based HetNet management and introduce ReaCritic, a reasoning transformer-based critic-model scaling scheme designed to address key challenges in HetNets.
4.1 DRL Markov Decision Process Formulation
DRL offers a principled framework for sequential decision-making by enabling an agent to interact with an environment $\mathcal{E}$. At each discrete time step $t$, the agent observes a state ${\bm{s}}_{t}\in\mathcal{S}$ from the state space and selects an action ${\bm{a}}_{t}\in\mathcal{A}$ according to its learned policy. The environment responds with a reward signal $r_{t}^{\mathcal{E}}\in\mathbb{R}$, and transitions to a new state ${\bm{s}}_{t+1}$ based on a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The learning objective is to optimize the expected cumulative reward by modeling the agent–environment interaction as an MDP.
@@ 4.1 DRL Markov Decision Process Formulation
    where $\tau_{m}\in\{0,1,\ldots,\tau_{\max}\}$ denotes the user type.Each user type corresponds to a distinct set of preferences over multiple objectives, encoded by weights $\alpha_{m}^{(i)}$ $\left(i=1,\ldots,5\right)$ in the utility function $Q_{m}$. These profiles are hidden from the network manager, introducing uncertainty into the learning process and increasing the complexity of policy adaptation.
@@ 4.2 ReaCritic Design
ReaCritic introduces a two-dimensional reasoning architecture designed to enhance the expressiveness and generalization capability of the critic network. Unlike conventional single-stream Transformer critics, our architecture is characterized by a structural modification that instantiates parallel reasoning paths. This design allows the framework to transcend static function approximation, which is otherwise limited to the (1,1) configuration where $H=1$ and $V=1$ represents a standard Transformer-based baseline. The first stage, denoted as Horizontal Reasoning (HRea), performs token-level expansion by generating multiple parallel state-action embeddings to simulate diverse perspectives. This mechanism is structurally distinct from standard ensemble methods; rather than training multiple independent networks, HRea expands the representational breadth within a single forward pass by constructing parallel attention branches that process replicated and diversified tokens simultaneously. The second stage, denoted as Vertical Reasoning (VRea), applies stacked Transformer blocks to process these tokens hierarchically. This vertical stacking facilitates hierarchical logic distillation across the parallel branches generated by HRea, enabling the model to refine complex dependencies through recursive abstraction rather than simple layer stacking.
As illustrated in Fig. 3, the HRea and VRea steps together enable ReaCritic to support fine-grained relational modeling and robust Q-value estimation in complex, high-dimensional HetNets. The whole procedure is provided in Algorithm 1. The architectural blueprint of ReaCritic is predicated on the Transformer mechanism to encapsulate the complex global dependencies inherent in highly coupled HetNets. Unlike convolutional or recurrent structures, which are often constrained by localized receptive fields or sequential processing bottlenecks, the self-attention mechanism facilitates a permutation-invariant modeling of all-to-all correlations among heterogeneous users. This structural property is essential for representing the intricate interference regimes where the operational state of a single mobile terminal significantly influences the global network utility. Consequently, the adoption of a transformer-based backbone is a principled choice to ensure that the Critic network possesses the necessary structural capacity to approximate the high-dimensional value functions of modern wireless environments.
Furthermore, the synergy between horizontal and vertical reasoning represents a systematic orchestration of inference-time compute rather than an empirical aggregation of disparate modules. The HRea module is designed to expand the representational breadth through parallel tokens to explore diverse state-action correlations, whereas the VRea module enables the hierarchical distillation of logical abstractions across stacked layers. This dual-axis scaling methodology ensures that the computational effort of the Critic is adaptively aligned with the intrinsic complexity of the task, thereby providing a robust and theoretically grounded pathway for achieving value estimation stability. By shifting the focus from fixed-capacity modeling to dynamic reasoning-time scaling, the proposed architecture transcends the limitations of static empirical stacking and offers a scalable solution for resource management in non-stationary 5G/6G scenarios.
@@
Input
  • Environment Markov environment $\mathcal{E}$, replay buffer $\mathcal{D}$
  • DRL Algorithm $\mathcal{A}$ discount factor $\gamma$, actor network $\pi_{\theta}$ (if have), target networks $Q_{\bar{\phi}}$, $\pi_{\bar{\theta}}$, learning rates $\eta_{Q},\eta_{\pi}$, mini-batch size $M$
  • ReaCritic critic network $Q_{\phi}$ with parameters $\phi$, HRea steps $H$, VRea steps $V$
@@
6:   Sample action $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ according to DRL algorithm $\cal A$ policy rules
7:   Execute $a_{t}$, observe $r_{t}$, $s_{t+1}$, and termination flag $d_{t}\in{0,1}$
@@
11:   Update phase
12:  Sample mini-batch $(s_{t},a_{t},r_{t},s_{t+1})$} from $\mathcal{D}$
13:  (Module 1. Horizontal Reasoning Module)
14:  Compute state-action embedding $\mathbf{z}_{0}$ using (18)
15:  Instantiate parallel paths by expanding into $H$ tokens using (19) and add exploration noise via (20)
16:  (Module 2. Vertical Reasoning Module)
17:  Perform hierarchical logic distillation through $V$ transformer blocks using (21)
18:  Synthesize representations via step-wise attention (22) and predict final Q-value using (23)
19:  (Critic Loss and Parameter Update)
20:  Compute target Q-value $y_{t}$
@@
21:  Compute critic loss $\mathcal{L}Q=\frac{1}{M}\sum{i=1}^{M}(Q_{\phi}(s_{t},a_{t})-y_{t})^{2}$
22:  Update critic parameters $\phi\leftarrow\phi-\eta_{Q}\nabla_{\phi}\mathcal{L}_{Q}$
23:  if actor $\pi_{\theta}$ exists then
24:   Compute policy gradient based on the reasoned Q-estimates using (25)
25:   Update actor  $\theta\leftarrow\theta+\eta_{\pi}\nabla_{\theta}J$
26:  end if
27:end for
Output Trained DRL algorithm $\cal A$
4.2.1 Horizontal Reasoning Path
@@
However, unlike LLMs that sequentially generate tokens, DRL critic networks operate under a single-shot evaluation paradigm. Therefore, to mimic multi-step reasoning, ReaCritic employs a horizontal reasoning expansion strategy whereby the model replicates the input state-action token multiple times and augments each copy with a unique positional embedding, effectively simulating parallel reasoning paths.
@@
where $\mathbf{W}{\text{embed}}\in\mathbb{R}^{(d{\bm{s}}+d_{\bm{a}})\times d_{\rm h}}$ is the trainable projection matrix, and $\mu(\cdot)$ denotes layer normalization. To enable multi-path reasoning, the base embedding is expanded into a sequence of $H$ horizontally parallel reasoning tokens as
@@
This horizontally extended token sequence is then passed to the subsequent vertical reasoning transformer layers, where latent value estimation is refined through attention-based abstraction. Intuitively, the expansion mechanism defined in Equations 19 through 21 simulates the generation of parallel thought paths, where each reasoning token explores a unique representational subspace of the state-action pair to mitigate the risks of localized feature bias in high-dimensional HetNet environments. The key intuition behind HRea is to provide the critic network with multiple perspectives on the same decision point, enhancing robustness, exploration, and its ability to model high-order interactions in complex environments. This mechanism systematically broadens the attention span of the Critic, allowing it to evaluate multiple potential interference patterns simultaneously rather than relying on a collapsed mean-field approximation typically found in standard architectures.
4.2.2 Vertical Reasoning Path
@@
This design is motivated by the idea of latent multi-step reasoning in LLMs [13], where deep transformer layers repeatedly refine representations in hidden space to support complex, context-dependent computation. Here, VRea does not explicitly produce symbolic CoT but instead performs recursive abstraction within the vector space. This allows the critic to simulate structured inference trajectories by reprocessing the latent input across multiple transformer layers, enabling deeper exploration and more robust Q-value modeling. Unlike simple feed-forward stacking found in standard models, this recursive distillation process ensures that the architectural depth translates directly into logical expressiveness across the parallel branches, effectively bridging the gap between high-level reasoning and low-level value estimation. Additionally, VRea also shares conceptual links with diffusion policy [6] and iterative policy refinement, where each step in the stack can be interpreted as an implicit planning operation conditioned on both prior layers and current observations. Compared with simple feed-forward architectures, the VRea stack increases reasoning depth without increasing parameter count proportionally, and facilitates richer representation learning through hierarchical feature composition.
@@
where $\Psi^{(v)}(\cdot)$ denotes the $v_{\rm th}$ transformer block. Each layer refines the horizontal token sequence by integrating long-range dependencies and capturing high-order interactions through attention and non-linear transformation. Specifically, while HRea focuses on representational breadth, the recursive operations in Equations 22 and 23 facilitate hierarchical logic distillation by iteratively filtering redundant information and accentuating critical dependencies necessary for precise value estimation. Additionally, after $V$ layers, the final representation $\mathbf{Z}^{(V)}\in\mathbb{R}^{B\times H\times d_{\rm h}}$ is aggregated using attention over the $H$ horizontal reasoning steps as
@@
This vertical reasoning mechanism enables ReaCritic to synthesize structured value estimates by abstracting state-action pairs, resulting in improved generalization and temporal stability across complex HetNet scenarios. To further ensure technical clarity, we have introduced explicit modular labels within Algorithm 1 to distinguish between the Horizontal and Vertical reasoning stages. Each module is now preceded by structural comments that delineate the input-output flow, providing a transparent roadmap for implementing the proposed inference-time scaling strategy in large-scale wireless resource management tasks.
4.2.3 Integration with DRL algorithms
Figure 4: Integration of ReaCritic into DRL algorithms. Part 1 illustrates the reasoning-augmented SAC pipeline. Parts 2 to 5 demonstrate compatibility with other actor–critic methods, including TD3, A3C, DDPG, and PPO.
As shown in Fig. 4, ReaCritic is designed as a plug-and-play critic module compatible with standard actor-critic DRL algorithms. While our implementation is based on the SAC framework [16], the design can be generalized to other value-based methods such as Twin Delayed Deep Deterministic (TD3) [12], Deep Deterministic Policy Gradient (DDPG) [25], and Asynchronous Advantage Actor-Critic (A3C) algorithms [3].
@@ 4.2.3 Integration with DRL algorithms
4.2.4 Theoretical Complexity and Convergence Analysis
@@
The per-update computational complexity of ReaCritic is $\mathcal{O}(BH^{2}d_{\rm h}+BHVd_{\rm h}^{2})$, where $B$ is the batch size, $H$ is the number of HRea steps, $V$ is the number of vertical Transformer blocks, and $d_{\rm h}$ is the hidden dimension. The first term corresponds to the cost of multi-head self-attention across $H$ tokens per layer, while the second term captures the complexity of feed-forward transformations in each Transformer block. This structure scales linearly with $B$ and $V$, and quadratically with $H$ and $d_{\rm h}$. Compared to standard MLP-based critics with complexity $\mathcal{O}(Bd_{\rm h}^{2})$, the added cost remains moderate for lightweight settings and can be efficiently batched on modern GPUs. The rationale for this increased complexity resides in the necessity of representational breadth, as large-scale HetNets require the structural capacity to explore vast state-action correlations that traditional MLPs cannot effectively capture. This overhead is mitigated by the architectural synergy with modern AI hardware, where SIMD (Single Instruction, Multiple Data) parallelization ensures that the increase in wall-clock time remains sub-linear relative to theoretical FLOPs. Consequently, ReaCritic favors scalability by processing multiple reasoning tokens simultaneously, justifying the marginal latency increase through significant gains in convergence stability.
Furthermore, the hyperparameters $H$ and $V$ offer a controllable trade-off between logical depth and representation space. While increasing the vertical depth $V$ scales computational cost linearly for enhanced abstraction, scaling the horizontal breadth $H$ expands the exploration of state-action patterns through parallel attention heads without a prohibitive penalty on sequential processing time. This structural flexibility allows network operators to dynamically adjust the reasoning-time compute of ReaCritic to match available hardware resources and the non-stationary demands of the HetNet environment, ensuring the framework remains practically tractable for large-scale deployments. More importantly, ReaCritic is designed to support scaling along both the depth ($V$) and breadth ($H$) dimensions, enabling deployment of reasoning-capable critics. This two-dimensional scaling path aligns with recent trends in inference-time compute augmentation and allows the model to serve both compact on-device learning and large-scale centralized training when compute resources permit.
@@
Regarding convergence, although ReaCritic increases model depth and nonlinearity, its use of pre-normalized residual blocks and standard Transformer modules facilitates stable training. Since it is integrated into conventional DRL update rules, e.g., Bellman backup and policy gradients, the theoretical convergence properties of the base algorithms, such as SAC or TD3, remain valid, assuming the critic maintains universal approximation capacity. The added reasoning depth improves the expressiveness of the value function, which can reduce approximation bias and accelerate convergence. Additionally, stepwise attention and horizontal perturbation act as implicit regularization, improving stability and credit assignment in complex, non-stationary environments.
From a deployment perspective, the feasibility of ReaCritic within 5G/6G ecosystems is governed by the critical trade-off between signaling overhead, coordination stability, and execution latency. First, regarding the communication footprint, the centralized training paradigm is essential to mitigate the non-stationarity inherent in coupled wireless environments. Although aggregating global state information incurs a signaling overhead during the learning phase, this investment is justified by the resulting stability gain, which enables the network to converge toward a robust joint resource allocation policy that suppresses mutual interference. Crucially, this centralization does not preclude real-time responsiveness. The architecture supports a decoupled execution model where the trained Actor is deployed for distributed inference at individual access points. This ensures that the sophisticated reasoning of the Critic guides the optimization process without imposing communication bottlenecks during the ultra-low latency transmission phase.
Second, regarding computational complexity, the overhead of reasoning-time scaling is systematically manageable through hardware-software synergy. Although the theoretical complexity scales quadratically with the horizontal breadth $H$, the multi-head attention mechanism is inherently amenable to parallelization. By leveraging dedicated edge AI accelerators or gNodeB-integrated tensor cores, the execution time for standard configurations can be maintained within the sub-millisecond thresholds of the 5G Transmission Time Interval (TTI). Furthermore, the structural flexibility of ReaCritic allows network operators to dynamically adjust the reasoning depth $V$ and breadth $H$ to match available hardware resources, ensuring that the framework remains both practically tractable and theoretically sound for large-scale deployments.
5 Numerical Results